OpenStudy (anonymous):

Exponential functions?? Help! F(2/3)=64 x the answer is 16? g(3/4)= 3(16) x I'm just really confused how they are solving for this? Thanks!

OpenStudy (anonymous):

f(2/3)=64(2/3) and g(3/4)=3(16)*(3/4) You need to plug the values in for x, and then multiply them out. Does that make sense, see what you get and we will make sure you are right!

OpenStudy (anonymous):

ok, i understand that part. but when it comes to solving for the root it gets confusing for number one I got 64 2/3 plugged in for x and number two 3(16)3/4

OpenStudy (anonymous):

You mean 64*(2/3), correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So to simplify that, you should get 128/3 which is simplest form. For the second one, 3*16=48. Take that 48 and multiply it by (3/4), so 48*(3/4)=144/4, which simplifies to 36.

OpenStudy (anonymous):

Ok great thanks alot!

OpenStudy (anonymous):

No worries, do you understand everything?

OpenStudy (anonymous):

Yes thanks!